# Learning an Adaptive and View-Invariant Vision Transformer for Real-Time UAV Tracking

### v2

@@ IV-B Comparison with Lightweight Trackers
4) Attribute-Based Evaluation: We also performed attribute-based evaluations to assess the robustness of our trackers (i.e., AVTrack-DeiT and AVTrack-MD-DeiT) in challenging scenarios involving viewpoint changes. This includes comparing our trackers with 16 SOTA lightweight trackers under challenges such as ‘Camera Motion’ on the DTB70 and VisDrone2018 and ‘Viewpoint Change’ on the UAV123, as shown in Fig. 4. Notably, the DTB70 and VisDrone2018 datasets do not have a dedicated subset for scenarios involving viewpoint changes. As an alternative, we used the ‘Camera Motion’ subset to evaluate performance in such scenarios, as it is a key factor contributing to viewpoint changes in visual tracking. For instance, when a camera or the object itself moves rapidly or changes direction suddenly, it can result in varying views of the target in the captured images. As observed, our tracker ranks first in ‘Camera Motion’ on VisDrone2018, and ‘Viewpoint Change’ on UAV123, exceeding the second-best tracker by 1.0%/0.5% and 1.0%/0.6% in precision/success rate, respectively, with only slight gaps of 0.1% and 0.6% to the top tracker in the ‘Camera Motion’ on DTB70. Therefore, this attribute-based evaluation validates, both directly and indirectly, the superiority and effectiveness of the proposed method in addressing the challenges associated with viewpoint change.
@@
Figure 5: The qualitative evaluations were performed on five video sequences from the following datasets: DTB70 [85], UAVDT [86], VisDrone2018 [87], UAV123@10fps [88], and UAV123 [88]. The evaluated sequences include Surfing12, S1607, uav0000180_00050_s, person10, and truck1.
@@
4) Qualitative evaluation: To intuitively showcase the tracking results in UAV tracking scenes, qualitative tracking results of our trackers (i.e., AVTrack-MD-DeiT and AVTrack-DeiT) alongside six SOTA UAV trackers are visualized in Fig. 5. Four video sequences, selected from different benchmarks and scenarios, are presented for demonstration: Surfing12 from DTB70, S1607 from UAVDT, uav0000180_00050_s from VisDrone2018, person10 from UAV123@10fps, and truck1 from UAV123. The upper right corner features a selectively zoomed and cropped view of the tracked objects within the corresponding frames for better visualization. As observed, compared to the other SOTA UAV trackers, our tracker tracks the target objects more accurately in all challenging scenes, including occlusion (i.e., S1607 and uav0000180_00050_s), low resolution (i.e., Surfing12), scale variations (i.e., in all sequences), and viewpoint change (i.e., in all sequences). In these cases, our method significantly outperforms others and offers a more visually appealing result, highlighting the effectiveness of the proposed approaches for UAV tracking.
@@
TABLE III: Precision (Prec.), success (Succ.), and average GPU speed (Avg.FPS) comparison between our AVTrack-Deit and DL-based tracker on five UAV tracking benchmarks.
@@ IV-C Comparison with Deep Trackers
To further validate the superiority of our trackers, we compare AVTrack-DeiT and AVTrack-MD-DeiT with 16 deep SOTA trackers, including DiMP [111], PrDiMP [112], TrSiam [113], TransT [114], AutoMatch [115], SparseTT [116], CSWinTT [117], SimTrack [13], OSTrack [12], ZoomTrack [16], SeqTrack [118], MAT [119], ROMTrack [120], DCPT [121], HIPTrack [19], and EVPTrack [18] ,across five UAV tracking datasets. The evaluation results are shown in Table III, which shows precision (Prec.), success rate (Succ.), their average (Avg.), and average GPU speed (Avg.FPS). As shown, our AVTrack-MD-DeiT and AVTrack-DeiT excel by achieving the fastest and second-fastest speeds while maintaining high performance comparable to the SOTA deep trackers. On average, although AVTrack-MD-DeiT and AVTrack-DeiT exhibit slight gaps in average precision and success rate compared to HIPTrack [19], EVPTrack [18], and ROMTrack [120], they are noticeably slower than our methods in terms of GPU speed. Specifically, our AVTrack-MD-DeiT is 9.7, 11.9, and 5.9 times faster than HIPTrack, EVPTrack, and ROMTrack, respectively. These results indicate that our method delivers both high precision and speed, validating its suitability for real-time UAV tracking that prioritizes efficiency as well as accuracy.
TABLE IV: Comparison of original teacher models and distillation models with varying teacher combinations: model complexity, performance, and average GPU speed. $\psi_{1}$, $\psi_{2}$, and $\psi_{3}$ represent the AVTrack-DeiT, AVTrack-ViT, and AVTrack-EVA, respectively.
@@ IV-D Ablation Study
Impact of the Number of Teachers: More teachers may provide more diverse knowledge. Table IV summarizes the model complexity, performance, and average GPU speed of the original teacher models and the distillation models using different teacher combinations. As shown, single-teacher distillation can significantly reduce model complexity during inference and achieve a notable speedup of 17%. However, it leads to a considerable decline in performance compared to the corresponding teacher model, with decreases of over 2.0% in average precision and success rate. In contrast, multi-teacher distillation methods, including double-teacher and triple-teacher distillation, not only reduce model complexity and provide a significant speedup but also result in a smaller performance decrease, with all drops remaining below 1.6% in average precision and success rate, and some even surpassing the performance of the teacher models. In the triple-teacher distillation, the student models (i.e., AVTrack-MD-ViT and AVTrack-MD-EVA) exhibit superior performance when compared to their corresponding teacher models with the same backbone. Specifically, AVTrack-MD-ViT achieves gains of 0.5% in average precision and 0.9% in success rate, while AVTrack-MD-EVA shows a 0.4% improvement in average precision, with only a 0.1% drop in success rate. Although AVTrack-MD-DeiT does not outperform AVTrack-DeiT, it experiences only a slight performance drop, with both average precision and success rate decreasing by less than 0.5%. Notably, in practice, the model complexity of AVTrack is higher than that of AVTrack-MD, as the activated blocks in AVTrack typically exceed half of the total blocks in the backbone. As a result, AVTrack is slower than AVTrack-MD when both use the same ViT backbone. These experimental results provide evidence for the effectiveness of our proposed multi-teacher knowledge distillation based on MI Maximization, with the performance gain attributed to the complementarity of multiple teacher models.
TABLE V: Performance of varying numbers of ViT blocks in student models across five UAV tracking datasets.
@@
7 83.7 64.6 82.4 60.5 84.8 63.8 83.7 66.3 84.1 66.3 83.7 64.3 5.8 1.7 287.7
8 82.9 64.0 82.8 60.8 85.3 64.4 84.1 66.6 84.7 66.7 84.0 64.5 6.2 1.8 263.5
TABLE VI: Ablation study on the weighting of loss $\mathcal{L}_{MI}$ in training AVTrack-MD-DeiT by varying $\eta$ from $0.1\times 10^{-1}$ to $0.1\times 10^{-6}$.
DTB70 UAVDT VisDrone UAV123 UAV123@10fps Avg.
$\eta$ Prec. Succ. Prec. Succ. Prec. Succ. Prec. Succ. Prec. Succ. Prec. Succ.
0 81.2 62.1 80.6 58.5 81.1 61.3 81.9 64.5 82.0 64.6 81.4 62.2
$0.1\times 10^{-1}$ 82.8 63.7 80.1 58.3 81.8 62.1 81.1 64.3 83.1 65.4 81.8 62.8
$0.1\times 10^{-2}$ 82.8 64.1 82.7 60.1 84.6 64.1 82.4 65.0 82.6 65.2 83.0 63.7
$0.1\times 10^{-3}$ 84.0 65.2 83.1 60.3 84.9 64.2 82.6 65.2 83.3 65.5 83.6 64.1
$0.1\times 10^{-4}$ 81.2 63.0 82.1 58.6 82.5 62.1 82.5 64.9 83.7 65.8 82.4 62.9
$0.1\times 10^{-5}$ 83.8 64.4 82.5 59.4 83.0 62.7 81.4 64.2 83.5 65.5 82.8 63.2
$0.1\times 10^{-6}$ 81.5 63.1 81.1 58.0 81.7 61.9 83.9 65.9 84.4 66.2 82.5 63.0
Impact of varying numbers of ViT blocks in student models: To closely investigate the impact of varying numbers of ViT blocks on performance in student models, we train the AVTrack-MD-DeiT student model using a range of block counts from 4 to 8. The evaluation results are shown in Table V. As observed, the number of ViT blocks in the student model directly affects both tracking performance and speed. On average, increasing the number of ViT blocks leads to an upward trend in accuracy and model complexity but a decline in tracking speed. As the number of blocks increases from 4 to 6, each additional block results in a greater than 1.0% improvement in both average precision and average success rate, an increase of more than 0.4M in parameters, and a rise of over 0.1G in FLOPs, accompanied by an approximately 10% decrease in speed. When the number of blocks exceeds 6, further increases do not lead to significant performance gains, while model complexity continues to rise substantially and speed continues to decrease significantly. Considering the balance between performance and speed, we have chosen to set the default number of ViT blocks in the student model to 6 in our implementation, which is half that of the teacher models.
Impact of Weighting the MI maximization-based multi-teacher knowledge distillation: To obtain the optimal weight $\eta$ for the proposed MI maximization-based multi-teacher knowledge distillation, we trained AVTrack-MD-DeiT using varied values of $\eta$ ranging from $0.1\times 10^{-1}$ to $0.1\times 10^{-6}$ with with a scale factor of 0.1. Evaluation results are shown in Table VI. As shown, our tracker achieves best performance when the loss weight ($\eta$) is set to $0.1\times 10^{-3}$. Additionally, we have observed that the second- and third-best performances across these datasets are scattered both above and below the value of $0.1\times 10^{-3}$, without any apparent patterns. This variation may be attributed to the inherent differences between these datasets. Setting a value of $0.1\times 10^{-1}$ for the loss weight ($\eta$) results in an average maximal difference of Avg. Prec. of 1.8% and a maximal difference of Avg. Succ. of 1.3%. These significant differences clearly demonstrate that the choice of weight a has a considerable impact on the tracking performance. To be more specific, when the proposed loss is appropriately weighted, it can enhance the tracking performance. However, if not properly weighted, it may have detrimental effects on the training of the tracking task.
@@
Method MD loss DTB70 UAVDT VisDrone UAV123 UAV123@10fps Avg. Avg.FPS
@@
310.6
Impact of Multi-Teacher Knowledge Distillation Based on MI Maximization: To demonstrate the superiority of the proposed multi-teacher knowledge distillation based on the MI maximization, we train the model using MSE loss and the proposed MI-based loss for multi-teacher knowledge distillation, respectively. The evaluation results across five datasets are presented in Table VII. As observed, using the MSE loss results in greater performance loss compared to our method. On average, employing the proposed $L_{MI}$ results in a notable enhancement over the MSE loss, with improvements of 1.9% in precision and 1.7% in success rate, respectively. Additionally, compared to the teacher model, our method incurs only a slight decrease of 0.5% in average precision and 0.3% in average success rate, while achieving a notable speedup of 20.9%. These comparisons show that the benefit of our multi-teacher knowledge distillation based on the MI maximization is that it can provide a more thorough measure of the relationship between features through MI-based loss, which enables the student model to learn a more effective representation from the teacher model. On the other hand, our MI-based loss is less sensitive to noise and outliers compared to MSE, making it particularly effective in noisy environments.
@@ IV-E Qualitative Results:
The proposed AVTrack enhances the robustness of Vision Transformers (ViTs) to viewpoint changes by enforcing invariance in the target’s feature representation. Building upon this, we present an improved tracker, dubbed AVTrack-MD, which integrates a novel multi-teacher knowledge distillation (MD) framework based on MI maximization. The extensive experimental results above show that the proposed MD framework not only greatly improves model efficiency but also enables the AVTrack-MD (i.e., the student models) to achieve performance comparable to, or even exceeding, that of AVTrack (i.e., the teacher models). To intuitively demonstrate the effectiveness of the proposed components, we visualize the feature maps of several examples in Fig. 6. These samples, generated by AVTrack-DeiT with and without the proposed distillation framework, feature targets undergoing viewpoint changes.
@@
Visualization of feature maps: In Fig. 6, the first row presents the original target images from UAV123@10fps [88], while the corresponding feature maps generated by AVTrack-DeiT*, AVTrack-DeiT, and AVTrack-MD-DeiT are displayed in the second, third, and fourth rows, respectively. Note that AVTrack-DeiT* refers to AVTrack-DeiT without the integration of the proposed VIR and AM components. As observed, the feature maps generated by AVTrack-DeiT and AVTrack-MD-DeiT demonstrate greater consistency with changes in viewpoint, while the feature maps from AVTrack-DeiT* exhibit more significant variations, particularly at different viewing angles. This suggests that the VIR component enhances the tracker’s ability to focus on targets experiencing changes in viewpoint, thereby improving overall tracking performance. Furthermore, the proposed MD allows AVTrack-MD-DeiT to inherit this capability from AVTrack-DeiT, enabling it to maintain similar robustness to the teacher models while also enhancing model efficiency. These qualitative results offer visual evidence of our method’s effectiveness in learning view-invariant feature representations using Vision Transformers (ViTs).
@@ IV-F Real-world Test
V Conclusions

### v3

@@ IV-B Comparison with Lightweight Trackers
5) Attribute-Based Evaluation: We also performed attribute-based evaluations to assess the robustness of our trackers (i.e., AVTrack-DeiT and AVTrack-MD-DeiT) in challenging scenarios involving viewpoint changes. This includes comparing our trackers with 16 SOTA lightweight trackers under challenges such as ‘Camera Motion’ on the DTB70 and VisDrone2018 and ‘Viewpoint Change’ on the UAV123, as shown in Fig. 5. Notably, the DTB70 and VisDrone2018 datasets do not have a dedicated subset for scenarios involving viewpoint changes. As an alternative, we used the ‘Camera Motion’ subset to evaluate performance in such scenarios, as it is a key factor contributing to viewpoint changes in visual tracking. For instance, when a camera or the object moves rapidly or changes direction suddenly, it can result in varying views of the target in captured images. As observed, our tracker ranks first in ‘Camera Motion’ on VisDrone2018, and ‘Viewpoint Change’ on UAV123, exceeding the second-best tracker by 1.0%/0.5% and 1.0%/0.6% in precision/success rate, respectively, with only slight gaps of 0.1% and 0.6% to the top tracker in the ‘Camera Motion’ on DTB70. Therefore, this attribute-based evaluation validates, both directly and indirectly, the superiority and effectiveness of the proposed method in addressing the challenges associated with viewpoint change.
@@
Figure 6: The qualitative evaluations were performed on three video sequences from the following datasets: UAVDT [85], VisDrone2018 [86], and UAV123@10fps [44]. The evaluated sequences include S1607, uav0000180_00050_s, and person10.
@@
4) Qualitative evaluation: To intuitively showcase the tracking results in UAV tracking scenes, qualitative tracking results of our trackers (i.e., AVTrack-MD-DeiT and AVTrack-DeiT) alongside six SOTA UAV trackers are visualized in Fig. 6. Three video sequences, selected from different benchmarks and scenarios, are presented for demonstration: S1607 from UAVDT, uav0000180_00050_s from VisDrone2018, and person10 from UAV123@10fps. The upper right corner features a selectively zoomed and cropped view of the tracked objects within the corresponding frames for better visualization. As observed, compared to the other SOTA UAV trackers, our tracker tracks the target objects more accurately in all challenging scenes, including occlusion (i.e., S1607 and uav0000180_00050_s), scale variations (i.e., in all sequences), and viewpoint change (i.e., in all sequences). In these cases, our method significantly outperforms others and offers a more visually appealing result, highlighting the effectiveness of the proposed approaches for UAV tracking.
@@
TABLE IV: Precision (Prec.), success (Succ.), and average GPU speed (Avg.FPS) comparison between our proposed trackers(i.e., AVTrack-DeiT and AVTrack-MD-DeiT) and DL-based tracker on five UAV tracking benchmarks.
@@ IV-C Comparison with Deep Trackers
To further validate the superiority of our trackers, we compare AVTrack-DeiT and AVTrack-MD-DeiT with 16 deep SOTA trackers across five UAV tracking datasets. The evaluation results are shown in Table IV, which shows precision (Prec.), success rate (Succ.), their average (Avg.), and average GPU speed (Avg.FPS). As shown, our AVTrack-MD-DeiT and AVTrack-DeiT excel by achieving the fastest and second-fastest speeds while maintaining high performance comparable to the SOTA deep trackers. On average, although AVTrack-MD-DeiT and AVTrack-DeiT exhibit slight gaps in average precision and success rate compared to HIPTrack [28], EVPTrack [27], and ROMTrack [118], they are noticeably slower than our methods in terms of GPU speed. Specifically, our AVTrack-MD-DeiT is 9.7, 11.9, and 5.9 times faster than HIPTrack, EVPTrack, and ROMTrack, respectively. These results indicate that our method delivers both high precision and speed, validating its suitability for real-time UAV tracking that prioritizes efficiency as well as accuracy.
TABLE V: Comparison of original teacher models and distillation models with varying teacher combinations: model complexity, performance, and average GPU speed. $\psi_{1}$, $\psi_{2}$, and $\psi_{3}$ represent the AVTrack-DeiT, AVTrack-ViT, and AVTrack-EVA, respectively.
@@ IV-D Ablation Study
Impact of the Number of Teachers: More teachers may provide more diverse knowledge. Table V summarizes the model complexity, performance, and average GPU speed of the original teacher models and the distillation models using different teacher combinations. As shown, single-teacher distillation can greatly reduce model complexity during inference, resulting in a significant speedup of over 17%. However, it leads to a considerable decline in performance compared to the corresponding teacher model, with decreases of over 2.0% in average precision and success rate. In contrast, multi-teacher distillation methods, including double-teacher and triple-teacher distillation, not only reduce model complexity and provide a significant speedup but also result in a smaller performance decrease, with all drops remaining below 1.6% in average precision and success rate, and some even surpassing the performance of the teacher models. In the triple-teacher distillation, the student models (i.e., AVTrack-MD-ViT and AVTrack-MD-EVA) exhibit superior performance when compared to their corresponding teacher models with the same backbone. Specifically, AVTrack-MD-ViT improves average precision by 0.5% and success rate by 0.9%, while AVTrack-MD-EVA shows a 0.4% improvement in average precision with only a 0.1% drop in success rate. Although AVTrack-MD-DeiT does not outperform AVTrack-DeiT, it experiences only a slight performance drop, with both average precision and success rate decreasing by less than 0.5%. Notably, in practice, the model complexity of AVTrack is higher than that of AVTrack-MD, as the activated blocks in AVTrack typically exceed half of the total blocks in the backbone. As a result, AVTrack is slower than AVTrack-MD when both use the same ViT backbone. These experimental results provide evidence for the effectiveness of our proposed multi-teacher knowledge distillation based on MI Maximization, with the performance gain attributed to the complementarity of multiple teacher models.
TABLE VI: Performance of varying numbers of ViT blocks in student models across five UAV tracking datasets.
@@
7 83.7 64.6 82.4 60.5 84.8 63.8 83.7 66.3 84.1 66.3 83.7 64.3 5.8 1.7 291.4
8 82.9 64.0 82.8 60.8 85.3 64.4 84.1 66.6 84.7 66.7 84.0 64.5 6.2 1.8 272.7
Impact of varying numbers of ViT blocks in student models: To closely investigate the impact of varying numbers of ViT blocks on performance in student models, we train the AVTrack-MD-DeiT student model using a range of block counts from 4 to 8. The evaluation results are shown in Table VI. As observed, the number of ViT blocks in the student model directly affects both tracking performance and speed. On average, increasing the number of ViT blocks leads to an upward trend in accuracy and model complexity but a decline in tracking speed. As the number of blocks increases from 4 to 6, each additional block results in a greater than 1.0% improvement in both average precision and average success rate, an increase of more than 0.4M in parameters, and a rise of over 0.1G in FLOPs, accompanied by an approximately 10% decrease in speed. When the number of blocks exceeds 6, further increases do not lead to significant performance gains, while model complexity continues to rise substantially and speed continues to decrease significantly. Considering the balance between performance and speed, we have chosen to set the default number of ViT blocks in the student model to 6 in our implementation, which is half that of the teacher models.
@@
Method $\mathcal{L}_{MD}$ DTB70 UAVDT VisDrone UAV123 UAV123@10fps Avg.
@@
Impact of Multi-Teacher Knowledge Distillation Based on MI Maximization: To demonstrate the advantage of our multi-teacher knowledge distillation through MI maximization, we train the model with both MSE loss and the proposed MI-based loss separately. The evaluation results across five datasets are presented in Table VII. As observed, employing MSE loss as the distillation loss results in a more significant performance decline compared to our approach. On average, employing the proposed $\mathcal{L}_{MD}$ results in a notable enhancement over the MSE loss, with improvements of 1.9% in precision and 1.7% in success rate, respectively. Additionally, compared to the teacher model, our method incurs only a slight decrease of 0.5% in average precision and 0.3% in average success rate. These comparisons show that the benefit of our multi-teacher knowledge distillation based on the MI maximization is that it can provide a more thorough measure of the relationship between features through MI-based loss, which enables the student model to learn a more effective representation from the teacher model. On the other hand, our MI-based loss is less sensitive to noise and outliers compared to MSE, making it particularly effective in noisy environments.
@@ IV-E Qualitative Results:
Visualization of feature maps: In Fig. 7, the first row presents the original target images from UAV123@10fps [44], while the corresponding feature maps generated by AVTrack-DeiT*, AVTrack-DeiT, and AVTrack-MD-DeiT are displayed in the second, third, and fourth rows, respectively. Note that AVTrack-DeiT* refers to AVTrack-DeiT without the integration of the proposed VIR and AM components. As observed, the feature maps generated by AVTrack-DeiT and AVTrack-MD-DeiT demonstrate greater consistency with changes in viewpoint, while the feature maps from AVTrack-DeiT* exhibit more significant variations, particularly at different viewing angles. This suggests that the VIR component enhances the tracker’s ability to focus on targets experiencing changes in viewpoint, thereby improving overall tracking performance. Furthermore, the proposed MD allows AVTrack-MD-DeiT to inherit this capability from AVTrack-DeiT, enabling it to maintain similar robustness to the teacher models while also enhancing model efficiency. These qualitative results offer visual evidence of our method’s effectiveness in learning view-invariant feature representations using ViTs.
@@ IV-F Real-world Test
Figure 9: Analysis of failure cases: AVTrack-DeiT struggles when the target is fully occluded, extremely small, or visually similar to surrounding objects, leading to tracking drift or loss.
V Limitations and future work
V-A Limitations
We present several representative failure cases in Fig. 9. Although our tracker demonstrates strong robustness to viewpoint changes by learning view-invariant representations, it lacks a mechanism for reconstructing or retrieving target features after full occlusion (as shown in the first row), limiting its ability to maintain temporal continuity. Additionally, its performance degrades when tracking extremely small, visually similar, or camouflaged targets (as shown in the last two rows), mainly due to limited feature resolution and insufficient contrast between the target and the background, which impairs accurate localization and increases the risk of drift or tracking failure.
V-B Future work
In future work, we plan to extend our framework by incorporating other advanced representation learning methods, such as contrastive learning and self-supervised feature alignment, to enhance robustness in the presence of full occlusion and background clutter. These techniques will help the tracker learn more discriminative and invariant features, improving its ability to differentiate the target from visually similar objects and maintain accurate tracking despite temporary loss of visual cues.
VI Conclusions